# ContraGen: A Multi-Agent Generation Framework for Enterprise Contradictions Detection

### v1

@@ III-B Contradiction Mining Agent
    Top-$k$ Semantic Filtering: For each source sentence, cosine similarity is computed against candidate targets using msmarco-distilbert-base-v3 [18], embeddings, fine-tuned for passage retrieval. The top-$k$ (k=5) pairs above threshold $\theta_{s}$ are retained, while short sentences (<10 words), numbers, and bullet points are filtered out. Duplicate pairs are removed via text-based hashing, effectively reducing the search space while preserving high-relevance candidates.
@@
    NLI-Based Contradiction Classification: The filtered sentence pairs are evaluated using a pretrained Natural Language Inference (NLI) model (facebook/bart-large-mnli) [19] to assign a label $\ell_{\text{NLI}}\in{\text{contradiction},\text{neutral},\text{entailment}}$ with confidence $p_{\text{NLI}}\in[0,1]$. Pairs labeled as contradictions or with $p_{\text{NLI}}\leq\theta_{\text{conf}}$ ($\theta_{\text{conf}}=0.7$) are forwarded for further assessment by a large language model (LLM).
@@
    LLM-Based Contradiction Judgment: For the retained pairs, GPT-4o [3] is prompted (see appendix A-C) to assess whether both statements can be simultaneously true. It returns a binary contradiction label $\ell_{\text{LLM}}\in{0,1}$, a justification, and a confidence score $p_{\text{LLM}}\in[0,1]$ reflecting its certainty.
@@
    where $\ell_{\text{NLI}},\ell_{\text{LLM}}\in\{0,1\}$ represent binary contradiction labels (1 = contradiction), and weights $w_{\text{NLI}},w_{\text{LLM}}$ are derived from the model confidences:
@@
Human-in-the-Loop (HITL) annotations are essential for ensuring both the realism of contradiction generation and the reliability of contradiction detection in long-form enterprise texts. Automated methods can inject and flag potential inconsistencies, but human expertise is required to validate plausibility and establish high-quality ground truth. For example, authority contradictions (e.g., Board approval vs. unilateral CEO approval) and cross-document conflicts (e.g., NDA permitting affiliate disclosure vs. master agreement prohibiting third-party disclosure) require expert interpretation to assess incompatibility. To quantify annotation reliability, we measured inter-annotator agreement (IAA) across both self-contradictions and pairwise contradictions. For self-contradictions ($n=141$), annotators achieved $95.0\%$ agreement with Cohen’s $\kappa=0.8973$ and Krippendorff’s $\alpha=0.8976$. For pairwise contradictions ($n=426$), the agreement was $96.7\%$ with Cohen’s $\kappa=0.8745$ and Krippendorff’s $\alpha=0.8746$. In practice, samples falling below a $90\%$ agreement threshold are adjudicated by a subject matter expert who selects the ground-truth label. Embedding such human judgment improves dataset fidelity, enhances model robustness, and ensures that contradiction detection remains interpretable and legally sound.
@@
III-C Retrieval Verifiability Agent
@@
This distinction makes contradiction evaluation diagnostic rather than merely descriptive. For example, consider two policy documents: “Employees are eligible for remote work after 30 days” vs. “Employees are eligible for remote work after 90 days.” If the official HR handbook specifies 90 days, the contradiction is retrieval-verifiable, indicating a retrieval weakness. If no authoritative source exists, it is retrieval-resistant, reflecting limitations in retrieval and the need for deeper reasoning. Labeling contradictions in this way localizes errors, making evaluation actionable for improving enterprise RAG systems.
IV Dataset Construction and Characteristics
To evaluate contradiction detection in a controlled yet realistic setting, a golden standard of 84 synthetic enterprise documents is constructed around a fictional multinational aerospace company which operates globally under strict confidentiality agreements: “a large aerospace company designing commercial aircraft and defense systems, with major manufacturing sites in Wichita, Toulouse, and Nagoya, and operating under strict international export controls.”
The corpus spans six high-level domains that reflect common organizational functions, each further divided into five subdomains:
  • HR and Personnel Management: HR Policy, Employee Relations, Talent Acquisition, Performance Management, Training and Development
  • Engineering and Development: Product Design, Software Development, Quality Assurance, R&D, Prototyping
  • Sales and Marketing: Lead Generation, Market Research, Campaign Management, Brand Strategy, Sales Enablement
  • Finance and Operations: Budgeting, Financial Reporting, Supply Chain, Procurement, Inventory Management
  • IT and Technical Support: Network Administration, Help Desk, Cybersecurity, System Integration, Cloud Services
  • Legal and Compliance: Contract Management, Regulatory Compliance, Litigation Support, Policy Drafting, Risk Assessment
After the base documents are generated using the above metadata and passing fluency checks, contradictions are systematically injected in a domain-specific manner. Each document in HR and Personnel Management, IT and Technical Support, and Legal and Compliance contains one self-contradiction, while an interleaving strategy within Engineering and Development, Finance and Operations, and Sales and Marketing modifies every second document to contradict the preceding one. This setup creates both intra- and inter-document conflicts, simulating realistic organizational scenarios. As shown in Table II, temporal and specificity dominate across contradiction types, with temporal inconsistencies prevalent in self-contradictions and specificity conflicts more common in pairwise cases. A single contradiction was counted under multiple categories if applicable as it can simultaneously be.
Although contradictions were injected by category, unintended conflicts (e.g., temporal inconsistencies) occasionally appeared. Human annotators tracked whether the injected contradictions were truly preserved as contradictions, or whether they blended smoothly into the text and no longer stood out. They also recorded whether these cases created detectable conflicts and whether the system correctly identified them, enabling a precise evaluation of robustness and sensitivity to subtle or spurious contradictions.
@@ IV Dataset Construction and Characteristics
Temporal 25 25
Numerical 7 6
Specificity 16 24
Policy Reversal 26 14
Authority 4 9
Process 10 11
Total 88 89
TABLE II: Distribution of contradiction types for self and pairwise contradictions.
V Results and Discussion
To evaluate contradiction detection in enterprise-scale documents, we implemented three complementary strategies: a Natural Language Inference (NLI) baseline, an LLM-based Judge, and a Hybrid model that combines the two. The NLI method offers scalability and precision but struggles to capture broader document context. In contrast, the LLM-based Judge provides richer semantic reasoning, yet remains vulnerable to prompt sensitivity, hallucination, and overconfidence in its outputs. The Hybrid approach balances these trade-offs by integrating NLI predictions with LLM judgments under a confidence-weighted scheme, yielding more stable and interpretable labels. Empirical results show that each detector identifies a partially overlapping subset of contradictions, underscoring the value of the Hybrid approach in achieving both breadth and reliability in contradiction mining. Since no single approach ensures an exhaustive coverage, a unified evaluation dataset was constructed:
@@ V Results and Discussion
This process resolved inconsistencies, consolidated duplicates, and produced a unified gold-standard dataset comprising 141 self-contradiction pairs and 426 pairwise-contradiction pairs, with representative examples shown in Tables III and IV.
Target Statement Contradiction Paragraph LLM Pred Human Label
A critical change is the implementation of a centralized digital recruitment platform across all sites. The implementation of the decentralized recruitment system at our Wichita, Toulouse, and Nagoya facilities will continue. FALSE TRUE
Following the initial assessment, documentation and evidence compilation is mandated to be completed within 30 calendar days. Documentation and evidence compilation must be completed within 90 calendar days. TRUE TRUE
Our compliance with international export controls has occasionally delayed shipments. Despite stringent export controls, these have been flawlessly integrated without costs. TRUE FALSE
A key finding is that components produced in Nagoya offer up to a 15% cost reduction compared to Wichita. As of June 18, 2024, our analysis indicates significant cost variations between Wichita and Nagoya. FALSE FALSE
@@
TABLE III: Examples of self-contradictions with LLM predictions vs. human labels.
Doc1 Chunk Doc2 Chunk LLM Pred Human Label
As of Feb 10, 2024, the wing design software implementation has been completed. The implementation of the wing design software has been delayed until March 20, 2024. TRUE TRUE
Conceptual Design Phase is initiated by senior aerodynamics engineer, Akira Tanaka. Conceptual Design Phase will involve avionics engineers, overseen by Dr. Aiko Takahashi. TRUE FALSE
All new aircraft models must undergo a five-phase quality assurance process. Prototyping for Next-Gen Avionics will be structured in four phases. FALSE TRUE
Quarterly audits will be conducted by Senior Compliance Analyst, Emma Dubois. Quarterly audits have been initiated by CFO Maria Fernandez. FALSE FALSE
@@
TABLE IV: Examples of pairwise contradictions with LLM predictions vs. human labels.
Method Accuracy Precision Recall F1
NLI 41.1% 40.0% 76.7% 52.6%
Hybrid 90.1% 92.6% 83.3% 87.7%
LLM Judge 82.3% 79.7% 78.3% 79.0%
TABLE V: Performance on self-contradiction detection.
Method Accuracy Precision Recall F1
NLI 29.3% 12.9% 63.1% 21.4%
Hybrid 87.8% 57.8% 73.8% 64.9%
LLM Judge 72.1% 34.1% 89.2% 49.4%
TABLE VI: Performance on pairwise contradiction detection.
The quantitative evaluation through precision, recall, and F1 metrics reveals consistent trends in detector performance:
1. Self-contradictions are easier to capture. As outlined in Table V both the hybrid and LLM judge methods achieve high accuracies ($\sim$82–90%) and F1-scores ($\sim$79–88%). In contrast, the pure NLI detector struggles, highlighting its limitations in handling nuanced business-process contradictions.
2. Pairwise contradictions are substantially more challenging. Table VI shows that NLI collapses to near-random performance (accuracy 41%, F1 53%), underscoring its inability to generalize across documents. The hybrid detector outperforms the LLM judge (F1 64.9% vs. 49.4%), driven by its higher precision (57.8%), though this comes at the cost of recall (73.8%). It provides a balanced trade-off, while the LLM judge catches most true contradictions but also flags many non-contradictions.
3. Distinct precision–recall trade-offs. The hybrid consistently favors precision and stability, while the LLM judge favors recall, surfacing a wider range of subtle cross-document conflicts albeit with more false positives. This divergence is particularly useful diagnostically: hybrid filters provide a reliable baseline, while LLM reasoning exposes contradictions that might otherwise be overlooked.
@@
The hybrid method achieves the best overall balance, reaching 87.8% accuracy and an F1 score of 64.9%, showing robustness in cross-document settings where subtle contextual shifts and entity references complicate detection. These results underscore that self-contradictions are easier to capture, whereas pairwise contradictions remain significantly harder, even for the hybrid model.
The strong performance gap between self and pairwise contradictions highlights an open challenge: while single-document consistency checks can be effectively addressed with current methods, multi-document contradiction detection requires more advanced techniques. Future work should explore approaches such as joint reasoning across documents or retrieval-aware prompting to bridge this gap. Strengthening this capability is particularly critical for enterprise applications like compliance monitoring, regulatory audits, and contract review, where surfacing subtle conflicts early can substantially reduce legal and operational risks.
VI Conclusion and Future Work
We present ContraGen, a framework for systematic contradiction generation and detection in long-form enterprise documents, moving beyond traditional sentence-pair benchmarks. The framework creates, annotates, and evaluates contradictions at the document level, enabling rigorous assessment of both intra-document and cross-document consistency. This design addresses a critical gap in enterprise AI, where contradictions frequently arise across business processes, policies, and regulatory documents. Empirical results indicate that LLM as a judge provides the most balanced trade-off between precision and recall for both self-and pairwise-contradictions, yet detecting cross-document contradictions remains a significant challenge due to contextual variability across heterogeneous sources.
Limitations. This work faces several limitations. (1) computational and API expenses scale directly with the volume of generated documents and injected contradictions, making large-scale experimentation costly; (2) the evaluation relies on synthetic yet realistic enterprise-style documents, since confidentiality and compliance constraints restrict access to proprietary corpora and large-scale public datasets are not yet available. (3) establishing reliable ground-truth labels continues to depend on human annotators, limiting full automation; (4) although the framework applies thresholds (e.g., perplexity caps, similarity filters, decision cutoffs), we did not systematically tune parameters, limiting reproducibility across computational settings. Finally, this study represents an initial step toward controllable contradiction generation and detection, laying the groundwork for future comprehensive LLM benchmarks.
Future Work. Future research will pursue several directions: (1) developing benchmark suites that integrate synthetic and semi-real documents (e.g., public policy manuals, contracts) to broaden generalizability; (2) enhancing cross-document contradiction detection through retrieval-aware prompting, multi-hop reasoning, and systematic evidence alignment across heterogeneous sources such as contracts, compliance reports, and regulatory filings; (3) conducting rigorous ablations and sensitivity analyses, including parameter tuning for similarity thresholds, NLI backbones, and LLM judges, to improve reproducibility; (4) evaluating the robustness of RAG pipelines when confronted with conflicting enterprise evidence, with attention to whether failures stem from retrieval gaps or reasoning gaps; and (5) considering practical guardrails for moving from synthetic to real-world corpora, including redaction, access control, and auditing mechanisms to ensure privacy and compliance. Together, these steps will help scale ContraGen from a controlled benchmark to a deployable evaluation framework for enterprise RAG systems.
@@ Appendix A Appendix
A-A Prompts for Data Generation

### v2

@@ 3.2 Contradiction Mining Agent
    Top-$k$ Semantic Filtering: For each source sentence, cosine similarity is computed against candidate targets using msmarco-distilbert-base-v3 (Reimers and Gurevych, 2019), embeddings, fine-tuned for passage retrieval. The top-$k$ (k=5) pairs above threshold $\theta_{s}$ are retained, while short sentences (<10 words), numbers, and bullet points are filtered out. Duplicate pairs are removed via text-based hashing, effectively reducing the search space while preserving high-relevance candidates.
@@
    NLI-Based Contradiction Classification: The filtered sentence pairs are evaluated using a pretrained Natural Language Inference (NLI) model (facebook/bart-large-mnli) (Yin et al., 2019) to assign a label $\ell_{\text{NLI}}\in{\text{contradiction},\text{neutral},\text{entailment}}$ with confidence $p_{\text{NLI}}\in[0,1]$. Pairs labeled as contradictions or with $p_{\text{NLI}}\leq\theta_{\text{conf}}$ ($\theta_{\text{conf}}=0.7$) are forwarded for further assessment by a large language model (LLM).
@@
    LLM-Based Contradiction Judgment: For the retained pairs, GPT-4o (OpenAI, 2024) is prompted (see appendix A.3) to assess whether both statements can be simultaneously true. It returns a binary contradiction label $\ell_{\text{LLM}}\in{0,1}$, a justification, and a confidence score $p_{\text{LLM}}\in[0,1]$ reflecting its certainty.
@@
    where $\ell_{\text{NLI}},\ell_{\text{LLM}}\in\{0,1\}$ represent binary contradiction labels (1-contradiction), and weights $w_{\text{NLI}},w_{\text{LLM}}$ are derived from the model confidences:
@@
Human-in-the-Loop (HITL) annotations are vital for ensuring both the realism of contradiction generation and the reliability of contradiction detection in long-form legal texts. Automated methods can flag inconsistencies, but human expertise is required to validate plausibility and establish reliable ground truth. To quantify annotation reliability, we measured inter-annotator agreement (IAA) across self- and pairwise contradictions: for self, agreement was $96.0\%$ with Cohen’s $\kappa=0.9143$, Krippendorff’s $\alpha=0.9146$; for pairwise, $94.77\%$ with $\kappa=0.828$, $\alpha=0.8278$. Samples below a $90\%$ threshold are adjudicated by a subject matter expert. Embedding such human judgment ensures dataset fidelity, model robustness, and legally sound contradiction detection.
@@
3.3 Retrieval Verifiability Agent
@@
This distinction makes contradiction evaluation diagnostic rather than descriptive. For example, consider two clauses: “Termination requires 30 days’ notice” vs. “Termination requires 90 days’ notice.” If a statute mandates 90 days, the contradiction is retrieval-verifiable, indicating a retrieval weakness. If no statute exists, it is retrieval-resistant, reflecting limitations in retrieval and the need for deeper reasoning. Labeling contradictions this way localizes errors, making evaluation actionable for improving legal RAG systems.
4 Dataset Construction and Characteristics
To evaluate contradiction detection in a controlled yet realistic setting, a golden standard of 50 synthetic legal documents is constructed around a fictional multinational company, Aerodyne Systems, which designs aerospace technologies and operates globally under strict confidentiality agreements. The corpus spans five legal domains—contract law, compliance and regulation, internal policy and governance, dispute resolution and litigation, and terms and service management—ensuring domain realism and structured variation across organizational functions.
Organization description - “A fictional multinational aerospace company called Aerodyne Systems that designs and manufactures proprietary aerospace technologies and operates across offices in the U.S., Europe, and Asia. It routinely engages in partnerships with external vendors, research institutes, and government agencies, requiring strict non-disclosure and licensing agreements.”
After passing fluency checks, contradictions are systematically injected in a domain-specific manner. Each document in Dispute Resolution and Litigation and Terms and Service Management contains one self-contradiction, while an interleaving strategy within Contract Law, Internal Policy and Governance, and Compliance and Regulation modifies every second document to contradict the preceding one. This setup creates both intra- and inter-document conflicts, simulating realistic organizational scenarios. As shown in Table 2, specificity and policy reversals dominate across contradiction types, with temporal inconsistencies prevalent in self-contradictions and specificity conflicts more common in pairwise cases, reflecting discrepancies in timelines, goals, and resource allocations.
@@ 4 Dataset Construction and Characteristics
Temporal 12 18
Numerical 5 1
Specificity 6 14
Policy Reversal 17 9
Authority 2 5
Process 6 9
Total 48 56
Table 2: Distribution of contradiction types for self and pairwise contradictions.
While contradictions were injected by category, unintended conflicts (e.g., temporal inconsistencies) occasionally arose. Human annotations tracked whether injected contradictions were successfully blended, created detectable conflicts, and were correctly identified, enabling precise evaluation of system robustness and sensitivity to spurious cases.
5 Results and Discussion
We benchmarked three detectors: NLI, LLM Judge, and Hybrid (NLI+LLM) on both self-contradictions and pairwise contradictions. As discussed in Section 3.2, NLI lacks contextual depth, while LLM Judge suffers from prompt sensitivity, limited evidence verification, hallucination, and overconfidence. The Hybrid approach integrates NLI predictions with LLM judgments using a confidence-weighted scoring strategy to produce a consolidated contradiction label. This combination was adopted to balance NLI’s scalability and precision with the LLM’s richer contextual reasoning capabilities, leading to more reliable contradiction detection. Contradictions were mined using three detectors, each identifying partially overlapping sets of pairs. Since no single detector achieved full coverage, a unified evaluation dataset was constructed:
@@ 5 Results and Discussion
This process resolved inconsistencies, consolidated duplicates, and produced a unified gold-standard dataset comprising 100 self-contradiction pairs and 306 pairwise-contradiction pairs, with representative examples shown in Tables 3 and 4.
Target Statement Contradiction Hybrid Pred Human Label
Effective from March 1, 2024, any disclosure of sensitive information without proper authorization will be considered a breach of this agreement. Unauthorized disclosure of sensitive information will not be considered a breach of this agreement until May 15, 2024. TRUE TRUE
Upon submission, summaries will automatically be distributed to the opposing party’s legal team and the appointed mediator, ensuring no delays in mediation proceedings. Summaries must be reviewed and approved by both the Lead Counsel and Senior Mediation Advisor, Priya Deshmukh, before submission. TRUE FALSE
It is critical to note that all joint development projects initiated after April 10, 2024, will adhere strictly to this framework. All joint development projects initiated after April 10, 2024, are exempt from adhering to this framework and will operate under an adaptive guideline model. FALSE TRUE
Seconded employees will retain full ownership of any intellectual property developed during their tenure with the joint venture. This agreement ensures that all creative outputs remain the exclusive property of the individual employee, irrespective of their parent company obligations. FALSE FALSE
@@
Table 3: Examples of self-contradictions with hybrid predictions vs. human labels.
Doc1 Target Statement Doc2 Contradiction Hybrid Pred Human Label
Aerodyne Systems is committed to maintaining full transparency and compliance, submitting all required reports by March 30, 2024, without exceptions. Aerodyne Systems will submit the required regulatory compliance reports by April 15, 2024, ensuring alignment with updated industry standards. This submission deadline accommodates necessary internal auditing processes. TRUE TRUE
It is imperative that all Aerodyne Systems employees and vendors understand that any breach of this agreement will result in immediate legal action and potential termination of contracts, as the Company maintains a zero-tolerance policy for violations. The vendor is authorized to disclose Confidential Information to third parties immediately upon the termination of this Agreement, without requiring prior written consent from Aerodyne Systems. TRUE FALSE
Upon termination, the Vendor agrees to return or destroy all documents containing or reflecting Confidential Information and to confirm completion of this process in writing. The vendor is authorized to disclose Confidential Information to third parties immediately upon the termination of this Agreement, without requiring prior written consent from Aerodyne Systems. FALSE TRUE
The Crisis Management Policy is interpreted and enforced by the Internal Policy and Governance team led by Fiona Zhang, Director of Internal Policy. The Internal Policy and Governance team, led by Dr. Ravi Kumar, Policy Compliance Officer, and supported by Mei Lin Tan, Legal Affairs Specialist, is responsible for overseeing the Conflict of Interest Declaration process. FALSE FALSE
@@
Table 4: Examples of pairwise contradictions with hybrid predictions vs. human labels.
The performance comparison of contradiction miners, summarized in Table 5, highlights clear differences across the three approaches. For self-contradiction detection, the NLI-only model achieves high recall (81.6%) but suffers from low precision (37.3%), frequently misclassifying non-contradictions as contradictions. The LLM-only model is more conservative, yielding higher precision (74.4%) but slightly lower recall (76.3%). The hybrid approach outperforms both, achieving 92.0% accuracy and an F1 score of 89.5%, demonstrating that combining NLI-based filtering with LLM judgment reduces false positives while maintaining strong recall.
@@
Pairwise contradiction detection proves substantially more challenging. The NLI-only model again exhibits high recall (66.7%) but extremely poor precision (16.0%), while the LLM-only model improves performance (F1 = 46.9%) but tends to over-predict contradictions. The hybrid method achieves the best overall balance, reaching 89.5% accuracy and an F1 score of 70.9%, showing robustness in cross-document settings where subtle contextual shifts and entity references complicate detection. These results underscore that self-contradictions are easier to capture, whereas pairwise contradictions remain significantly harder, even for the hybrid model.
Self-Contradiction Pairwise Contradiction
Model A P R F1 A P R F1
NLI-only 41.0 37.00 81.6 51.2 36.3 16.0 66.7 25.9
LLM-only 81.0 74.4 76.3 75.3 66.7 31.9 88.2 46.9
Hybrid 92.0 89.5 89.5 89.5 89.5 66.1 76.5 70.9
Table 5: Evaluation results for Contradiction detection (in %)
Key insights emerge from our findings. NLI models provide high recall but suffer from low precision, whereas LLMs offer better contextual reasoning at the cost of coverage. The hybrid approach integrates their strengths, achieving consistently superior precision and recall. While hybrid or LLM-based methods effectively handle single-document consistency, cross-document contradiction detection remains challenging due to contextual shifts and fragmented evidence. This highlights that we are far from solving cross-document contradiction, underscoring the need for robust, realistic benchmarks like ours that expose subtle, retrieval-induced conflicts in high-stakes domains. Addressing this gap will require retrieval-aware prompting or multi-hop reasoning over evidence.
6 Conclusion and Future Work
This work introduces a framework for systematic contradiction injection within long-form legal documents, extending beyond short sentence-pair benchmarks. By generating and labeling contradictions at the document scale, the framework enables rigorous evaluation of both intra-document and cross-document consistency, addressing a critical gap in legal AI. Results from our evaluation of contradiction detection methods show that hybrid NLI+LLM methods achieve the best balance of precision and recall, especially for self-contradictions, whereas cross-document contradictions remain challenging due to contextual variation.
Limitations. One of the main limitations is the high API cost that is proportional to the number of documents we generate and the contradictions we inject. The evaluation relies on synthetic but realistic documents rather than large-scale proprietary corpora, limiting validation on naturally occurring contradictions. Human annotation is required for reliable ground-truth labeling, preventing full automation. The focus is on demonstrating controllable contradiction generation rather than exhaustive comparisons of LLM performance.
Future Work. Future directions include: (1) extending the framework into standardized benchmark suites that combine synthetic and real-world contradictions to overcome the current reliance on a limited corpus and enable broader, more representative model evaluation; (2) improving cross-document contradiction detection through retrieval-aware prompting and multi-hop reasoning to align evidence across heterogeneous sources; (3) systematically testing the performance of different LLMs as context validators; and (4) evaluating the ability and robustness of RAG pipelines when presented with conflicting evidence to surface, resolve, or isolate contradictions during generation.
@@ Appendix A Appendix
A.1 Prompts for Data Generation